OpenStudy (anonymous):

WILL FAN AND/OR MEDAL what is the solution to this system, use substitution y=x+6 y=2x

Directrix (directrix):

You are looking for an ordered pair (x,y) should that the values of x and y cause both equations to be true statements at the same time. y=x+6 y=2x ------- Try substitution. y = x+ 6 But y also equals 2x. So, 2x = x + 6. Solve for x and post what you get. Then, we'll find the value of y. Okay? @kristyn-whitaker

OpenStudy (anonymous):

3x+6 i think

OpenStudy (anonymous):

@Directrix

Directrix (directrix):

Your task is to solve this equation for x: So, 2x = x + 6. 2x = x + 6 Subtract x from each side. So, 2x = x + 6. -x -x --------------- x = ? @kristyn-whitaker

OpenStudy (anonymous):

-1x=2 then right?

OpenStudy (anonymous):

@Directrix

Directrix (directrix):

No, you are losing the equality sign.

Directrix (directrix):

So, 2x = x + 6. ------------- Question: What is 2x - 1x equal to? @So, 2x = x + 6.

OpenStudy (anonymous):

im so confused they teach us so different here than what your trying to teach me

Directrix (directrix):

If you had two dogs and one of them ran away, how many dogs would you have left? This is related to the question 2x - 1x = ?. We are close to a solution. So, how many dogs are left?

OpenStudy (anonymous):

1

Directrix (directrix):

>>-1x=2 then right? This is not correct. ------------- Yes, one dog is left so 2x - 1x = x, Now, for the value of x. What is x + 6 - x =? In this scenario, you had one dog and 6 birds. The dog ran off and left you with the 6 birds. So x = 6.

Directrix (directrix):

Here's the algebra for that: Directrix Medals 0 You are looking for an ordered pair (x,y) should that the values of x and y cause both equations to be true statements at the same time. y=x+6 y=2x ------- Try substitution. y = x+ 6 But y also equals 2x. So, 2x = x + 6. -x = -x --------------- x = 6 ----> Now find y.

Directrix (directrix):

y=2x y = 2 * 6 = ? @kristyn-whitaker

OpenStudy (anonymous):

y=12?

Directrix (directrix):

Write the solution as an ordered pair. (6, ?) where ? is what you got for y. YES, y = 12 So, the answer is (6, 12) x = 6 and y = 12

OpenStudy (anonymous):

(6,12)

Directrix (directrix):

@kristyn-whitaker Question?

OpenStudy (mvpriest6):

i'm trying to send a formula but it wont send

OpenStudy (anonymous):

its cool but through all the confusion i got it right

Directrix (directrix):

Good to hear.

OpenStudy (mvpriest6):

okay well thanks for helping her @Directrix